OpenStudy (anonymous):

Solve the equation for x. 1/3(x+5)=k

OpenStudy (texaschic101):

1/3(x + 5) = k ----distribute through the parenthesis 1/3x + 5/3 = k --- subtract 5/3 from both sides 1/3x = k - 5/3 -- divide by 1/3 x = (k - 5/3)/(1/3) x = k/(1/3) - (5/3)/(1/3) x = k*3/1 - 5/3*3/1 x = 3k - 5

jhonyy9 (jhonyy9):

or easy way 1 --(x+5)= k cross multiplie 3 x+5=3k x=3k-5